# Bine Trees: Enhancing Collective Operations by Optimizing Communication Locality

### v1

@@ 1. Introduction
As high-performance computing (HPC) systems and data centers continue to grow (500, 2025; Atchley et al., 2023a; Meta, [n. d.]; Morgan, 2025), communication locality plays a crucial role in determining performance. In most HPC networks, communication with nearby nodes is generally faster than with distant nodes. This is particularly evident in scale-up networks, where bandwidth can be up to an order of magnitude higher than that of the scale-out network (De Sensi et al., 2024b; Hoefler et al., 2022). However, even when focusing solely on the scale-out network, locality remains important. While full-bandwidth scale-out networks such as fat trees can deliver high performance, they also come with significant cost and power consumption. As a result, oversubscribed (tapered) networks are often used. These are typically organized into fully connected groups, such as groups in Dragonfly/Dragonfly+ topologies (Kim et al., 2008; Shpiner et al., 2017; De Sensi et al., 2020; Atchley et al., 2023b) or subtrees in oversubscribed fat trees, which are then interconnected more sparsely through global links.
Reducing the traffic forwarded over global links is advantageous for several reasons. First, global links are often longer than local links, resulting in higher communication latency (De Sensi et al., 2020; Atchley et al., 2023b). Second, because global links are typically oversubscribed, they can sustain only a limited number of concurrent communications, often causing congestion when multiple ranks from the same or different applications contend for the same link, leading to higher latency and lower throughput (Chunduri et al., 2019; De Sensi et al., 2019; Smith et al., 2018; De Sensi et al., 2024b). This issue is amplified in oversubscribed low-diameter networks, where jobs are usually distributed across multiple groups. Such distribution helps arbitrary traffic patterns approximate uniform behavior, under which these topologies achieve high throughput (Wang et al., 2025; Yang et al., 2016; Kang et al., 2024; Blach et al., 2024; SchedMD, 2025b), but it also increases global link utilization. Third, global links consume more energy than local links, making it crucial to reduce their usage (Abts et al., 2010). Last, as the physical size of data centers and supercomputers approaches practical limits, tightly coupled applications run across multiple geographically distributed data centers (Luo et al., 2024; Gemini Team, 2023; OpenAI, 2025; Dylan Patel, Daniel Nishball and Jeremie Eliahou Ontiveros, [n. d.]). In such scenarios, communication between distinct data centers or supercomputers must traverse even slower Wide-Area Network (WAN) links, and improving communication locality is critical.
@@
For instance, consider the scenario in Fig. 1, where eight nodes are connected through a 2:1 oversubscribed fat-tree. Each leaf switch has two downlinks and one uplink, serving two nodes per switch. In a distance-doubling binomial tree (as in Open MPI (MPI, 2025a, b), top part of Fig. 1), the distance between communicating ranks doubles at each step (e.g., rank 0 first sends its data to rank 1, then to rank 2). In contrast, in a distance-halving tree (as in MPICH (MPICH, 2025a), bottom part of the Fig. 1), the distance halves at each step (e.g., rank 0 first sends its data to rank 4, then to rank 2). In both cases, each rank that has already received the $n$ bytes of the vector forwards them to another rank. The distance-doubling broadcast forwards a total of $6n$ bytes over global links ($2n$ at step $1$ and $4n$ at step $2$), compared to only $3n$ bytes for the distance-halving variant.
Figure 1. Traffic on global links for a broadcast collective using distance-doubling and distance-halving binomial trees.
[…]
Traffic on global links for a broadcast collective using distance-doubling and distance-halving binomial trees.
@@
Moreover, although ring algorithms reduce global link traffic compared to binomial trees and butterflies, their performance is suboptimal for large numbers of ranks and/or small to medium vector sizes (De Sensi et al., 2024a). Also, it is important to note that existing hierarchical and locality-aware collective algorithms may not always be effective. These algorithms typically require prior knowledge of the number of processes per group and often assume that jobs run with the same number of processes in each group (Bienz et al., 2022; Träff, 2006; Karonis et al., 2000; Kandalla et al., 2009; Graham et al., 2011). While such assumptions are reasonable for the number of processes per node, they may not hold when considering the partitioning of processes across subtrees or Dragonfly groups. In fact, the scheduler’s process-to-node allocation is not known in advance and is unlikely to result in an even distribution of processes across groups.
@@ 2.1. Binomial Trees Construction
A distance-halving binomial tree of order 0 consists of a single node. A tree of order $k$ is built recursively by linking two trees of order k1k-1, making one the child of the other. This structure is shown in the top part of Fig. 2, where each subtree is itself a binomial tree. When used in collective operations, each node maps to a rank, and each edge represents communication between ranks. Any mapping of ranks to nodes yields a valid algorithm, also for collectives involving computation (e.g., reduce), as long as the operator is associative, which MPI assumes by default.
Figure 2. Distance-halving binomial tree construction.
[…]
Distance-halving binomial tree construction.
@@ 2.2. Bine Trees Construction: Visual Intuition
The key idea behind Bine trees is to build a binomial tree by arranging ranks so that the distance between communicating ranks is smaller than in standard binomial trees. Shorter distances increase the likelihood that communication occurs within the same group (or within the same node), thereby benefiting from higher bandwidth and lower latency. Since our goal is to design an algorithm as simple and generic as standard binomial tree algorithms, and not tied to any specific topology, we do not consider the actual physical distance between nodes (e.g., by retrieving their location (Ma et al., 2011)).
In standard binomial trees, the distance between communicating ranks is defined by the difference in their identifiers. For instance, in distance-halving binomial trees, rank $6$ first communicates with rank $4$ (distance $2$) and then with rank $7$ (distance $1$). This provides a reasonable approximation of physical distance when ranks are mapped linearly across the system, which is often the default choice. For example, this is the case in Slurm’s block distribution (SchedMD, 2025a).
@@
We make the same assumption that ranks are mapped linearly. If this is not the case, a remapping to a block distribution can be applied when the communicator is created, and the remapped rank identifiers can be used instead of the original ones. For instance, on the systems used in this paper, hostnames are numbered consecutively across groups or subtrees, so it is enough to sort ranks by hostname. However, differently from binomial trees, we consider the modulo distance between rank identifiers. Namely, we arrange the $p$ ranks in a circle $0,1,\ldots,p-1$, and consider the distance between two ranks $r$ and $q$ as the minimum distance along the circle. I.e., $d(r,q)=\min((r-q)\bmod p,\,(q-r)\bmod p)$.
@@
We construct a distance-halving Bine tree recursively. A two-node Bine tree (Fig. 3, order 1) is identical to a standard binomial tree. To build a four-node (i.e., order 2) Bine tree, we connect an order 1 Bine tree (Fig. 3, [Uncaptioned image]), to a mirrored order 1 Bine tree (i.e., reflected on the vertical axis) and placed to its left (Fig. 3, [Uncaptioned image]). The mirrored tree is placed to the left rather than to the right to keep a shorter distance between the roots of the subtrees (one rather than two as in a binomial tree – see Fig. 2, [Uncaptioned image]). Ranks are numbered left to right starting from the root, wrapping around once the rightmost node of the tree is reached (e.g., on the 4-node tree, ranks $2$ and $3$ are those on the left of the root).
Figure 3. Distance-halving Bine trees construction.
[…]
Distance-halving Bine trees construction.
@@
The process generalizes recursively: an eight-node Bine tree is constructed by combining an order-2 Bine tree (Fig. 3, [Uncaptioned image]) and a mirrored order-2 Bine tree (Fig. 3, [Uncaptioned image]), placed side by side and arranged so that the modulo distance between the roots is as small as possible. In this case, placing the new four-node subtree on the right (rather than on the left) yields a root-to-root modulo distance of three, compared to five in the alternative placement. For comparison, the corresponding binomial tree would result in a distance of four (Fig. 2, [Uncaptioned image]). If $p$ is not a power of two, standard techniques used in binomial trees are applied (e.g., forwarding from excess ranks (MPICH, 2025a)). Moreover, if the root is rank $t\neq 0$, we apply a logical rotation by subtracting $t$ from all rank identifiers modulo $p$.
@@ 2.3.2. Determining the Communication Partner
In the broadcast, each rank needs to determine when to join the tree, or in other words, at which step it will receive data from its parent. We denote by $u$ the number of consecutive least significant bits that are equal to each other, starting from the least significant bit. For instance, for a 16-node Bine tree, $u=3$ for $1000$, and $u=2$ for $1011$. In a distance-halving Bine tree broadcast, rank $r$ receives data at step i=sui=s-u (with $i\in[0,s-1]$). As illustrated in Fig. 4 [Uncaptioned image], rank $8$ receives the data from its parent at step $i=1$, since $rank2nb(8)=1000$, and thus $u=3$, giving i=su=43=1i=s-u=4-3=1.
Figure 4. A 16-node (order 4) distance-halving Bine tree.
[…]
A 16-node (order 4) distance-halving Bine tree.
@@
with $\oplus$ denoting the XOR operation. In other words, at step $i$, each rank $r$ that has already received the data will send it to a rank $q$ whose negabinary representation differs in the least significant sis-i bits. For example, as shown in Fig. 4 [Uncaptioned image], at step $i=2$, rank $8$ sends data at rank $7$ since their negabinary representations ($1000$ and $1011$, respectively), differ for the least significant si=42s-i=4-2 bits. For clarity, the complete MPI code of the distance-halving Bine tree algorithm for the broadcast is provided in Listing 1. Notably, the implementation is no more complex than that of standard binomial trees (MPICH, 2025a; MPI, 2025a). For space reasons, we omit handling cases where the number of nodes is not a power of two; these can be managed similarly to standard binomial tree algorithms.
@@ 2.4.1. Theoretical Bound Computation
We can estimate the reduction in the distance between communicating ranks by considering the ratio:
@@
This implies that in Bine trees, communicating ranks are at a $\sim 33\%$ shorter distance compared to standard binomial trees. As a result, the probability that two communicating ranks are in different groups and thus need to communicate through global links is reduced. Quantifying the exact reduction in inter-group traffic is challenging, as it heavily depends on the specific topology and on how ranks are allocated to groups. However, we can estimate an upper bound by considering a worst-case scenario in which each group only has a single outgoing link to another group.
Let us focus on a specific step $i$ and assume the group size equals $\delta_{binomial}(i)$, such that all communications in the standard binomial tree cross group boundaries. In this worst-case scenario, all $\delta_{binomial}(i)$ communications would be directed to nodes in other groups, whereas for Bine trees this would only happen for $\delta_{bine}(i)$ communications. Therefore, the maximum reduction in global link traffic corresponds to the ratio $\delta_{bine}(i)/\delta_{binomial}(i)$, which, as shown in Eq. 2, is equal to $2/3$. In other words, Bine reduces by at most $33\%$ the traffic on global links.
@@ 2.4.2. Empirical Analysis
We additionally validated our assumptions empirically by analyzing one-week job allocation data from the Leonardo (Turisini et al., 2024) supercomputer and two-weeks data from the LUMI (EuroHPC Joint Undertaking, [n. d.]) supercomputer. Using Slurm commands (squeue and scontrol), we obtained the list of jobs running on the system and, for each job, the hostnames of the allocated nodes. We then identified the Dragonfly/Dragonfly+ group to which each node was assigned. On Leonardo, we relied on existing hostname-to-group mappings (CINECA, 2025), while on LUMI we extracted this information using the sinfo -N -o "%N %f" command. With this data, we determined the group assignment of each rank for every job. Based on this mapping we estimated, for each job, the global traffic of an allreduce operation using standard binomial and Bine trees, and computed the resulting traffic reduction achieved by Bine across allocations.
Figure 5. Reduction in global traffic on Leonardo and LUMI when using Bine trees compared to standard binomial trees.
[…]
Reduction in global traffic on Leonardo and LUMI when using Bine trees compared to standard binomial trees.
We present this analysis in Fig. 5, where we report the distribution of global traffic reduction across jobs using boxplots, grouped by the number of nodes. The data includes a total of $1116$ jobs on Leonardo and $1914$ jobs on LUMI. The largest observed jobs spanned up to $256$ nodes on Leonardo and $2048$ nodes on LUMI, matching the maximum job sizes allowed on the two systems. For clarity, outliers are not shown in the figure. No outliers above the theoretical upper bound of 33% were observed, coherently with our model (Eq. 2). A few outliers reached values as low as $200\%$, all occurring in small allocations (fewer than 64 nodes). As discussed in Sec.2.2, this happens because Bine minimizes distance modulo $p$ rather than the actual distance. In a few small allocations, this can increase global traffic compared to binomial trees, although on average, they remain comparable. In contrast, for larger allocations, Bine consistently outperforms standard binomial trees. The reduction in global link traffic increases with the number of nodes, as larger jobs typically span more groups, creating more opportunities for Bine to reduce inter-group communication.
@@ 3.1. Bine Butterflies
In a butterfly communication pattern, each rank exchanges data with another rank at every step. A Bine distance-halving butterfly can be seen as the superposition of multiple distance-halving Bine trees, each rooted at a different rank. The left part of figure 6 illustrates this view, highlighting the trees rooted at ranks 2 and 3 with solid orange and dotted red lines, respectively. To guarantee correctness, these trees must be arranged so that the send and receive decisions are consistent across all ranks.
Figure 6. An 8-node distance-halving Bine butterfly (left), and distance-doubling Bine butterfly (right).
[…]
An 8-node distance-halving Bine butterfly (left), and distance-doubling Bine butterfly (right).
We take as a reference the distance-halving tree rooted in rank 0. As discussed in Sec. 2.3, at step $i$ rank 0 communicates with a rank whose negabinary representation has the least significant sis-i bits set to 1, and the remaining bits set to 0, i.e., with rank:
@@
Trees rooted in even ranks $r$ have the same structure as the tree rooted in 0, but rotated to the right by $r$ positions. Moreover, since each rank communicates with another rank whose negabinary representation differs in the last $i+1$ bits, even ranks always communicate with odd ranks, and vice versa. Therefore, the tree rooted at an odd rank $r$ must be mirrored with respect to the tree rooted at the even rank r1r-1. Thus, we can determine the communication partner $q$ for any rank $r$ at step $i$ as:
@@
Since the distance between communicating ranks in Bine butterflies matches that of Bine trees, distance-halving Bine butterflies reduce global traffic by up to $33\%$ compared to standard distance-halving butterflies. Distance-halving butterflies are advantageous for collectives that increase the size of the exchanged data at each step, such as allgather, since in the last steps, where more data is transmitted, communication happens between closer nodes. Alternatively, it is possible to build a distance-doubling Bine butterfly by inverting the communication order, as shown in the right part of Fig. 6. This is preferable for collectives that exchange more data in the first steps, such as reduce-scatter (Sec. 4.3).

### v2

@@ 1. Introduction
As high-performance computing (HPC) systems and data centers continue to grow (top500; 10.1145/3581784.3607089; metacluster; colossus), communication locality plays a crucial role in determining performance. In most HPC networks, communication with nearby nodes is generally faster than with distant nodes. This is particularly evident in scale-up networks, where bandwidth can be up to an order of magnitude higher than that of the scale-out network (gpugpuinterconnect; hxmesh). However, even when focusing solely on the scale-out network, locality remains important. While full-bandwidth scale-out networks such as fat trees can deliver high performance, they also come with significant cost and power consumption. As a result, oversubscribed (tapered) networks are often used. These are typically organized into fully connected groups, such as groups in Dragonfly/Dragonfly+ topologies (dragonfly; dragonflyplus; sc2020; frontier) or subtrees in oversubscribed fat trees, which are then interconnected more sparsely through global links.
Reducing the traffic forwarded over global links is advantageous for several reasons. First, global links are often longer than local links, resulting in higher communication latency (sc2020; frontier). Second, because global links are typically oversubscribed, they can sustain only a limited number of concurrent communications, often causing congestion when multiple ranks from the same or different applications contend for the same link, leading to higher latency and lower throughput (gpcnet; sc2019; staci2018; gpugpuinterconnect). This issue is amplified in oversubscribed low-diameter networks, where jobs are usually distributed across multiple groups. Such distribution helps arbitrary traffic patterns approximate uniform behavior, under which these topologies achieve high throughput (randompl1; randompl2; randompl3; slimfly; slurm_spread), but it also increases global link utilization. Third, global links consume more energy than local links, making it crucial to reduce their usage (36462). Last, as the physical size of data centers and supercomputers approaches practical limits, tightly coupled applications run across multiple geographically distributed data centers (uno-tom; luo2024efficient; gemini_google; openai_yt; multidc). In such scenarios, communication between distinct data centers or supercomputers must traverse even slower Wide-Area Network (WAN) links, and improving communication locality is critical.
@@
For instance, consider the scenario in Fig. 1, where eight nodes are connected through a 2:1 oversubscribed fat-tree. Each leaf switch has two downlinks and one uplink, serving two nodes per switch. In a distance-doubling binomial tree (as in Open MPI (ompi-bcast-binomial; ompi-binomial), top part of Fig. 1), the distance between communicating ranks doubles at each step (e.g., rank 0 first sends its data to rank 1, then to rank 2). In contrast, in a distance-halving tree (as in MPICH (mpich-bcast-binomial), bottom part of the Fig. 1), the distance halves at each step (e.g., rank 0 first sends its data to rank 4, then to rank 2). In both cases, each rank that has already received the $n$ bytes of the vector forwards them to another rank. The distance-doubling broadcast forwards a total of $6n$ bytes over global links ($2n$ at step $1$ and $4n$ at step $2$), compared to only $3n$ bytes for the distance-halving variant.
Figure 1. Traffic on global links for a broadcast collective using distance-doubling and distance-halving binomial trees.:
@@
Moreover, although ring algorithms reduce global link traffic compared to binomial trees and butterflies, their performance is suboptimal for large numbers of ranks and/or small to medium vector sizes (swing). Also, it is important to note that existing hierarchical and locality-aware collective algorithms may not always be effective. These algorithms typically require prior knowledge of the number of processes per group and often assume that jobs run with the same number of processes in each group (10.1145/3555819.3555825; 10.1007/11846802_16; 846009; 5160896; 10.1109/CCGrid.2011.42). While such assumptions are reasonable for the number of processes per node, they may not hold when considering the partitioning of processes across subtrees or Dragonfly groups. In fact, the scheduler’s process-to-node allocation is not known in advance and is unlikely to result in an even distribution of processes across groups.
@@ 2.1. Binomial Trees Construction
As shown in Fig. 2, a distance-halving binomial tree of order 0 consists of a single node. A tree of order $k$ is built recursively by linking two trees of order k1k-1, making one the child of the other. When used in collective operations, each node maps to a rank, and each edge represents communication between ranks. Any mapping of ranks to nodes yields a valid algorithm, also for collectives involving computation (e.g., reduce), as long as the operator is associative, which MPI assumes by default.
Figure 2. Distance-halving binomial tree construction.:
@@ 2.2. Bine Trees Construction: Visual Intuition
The key idea of Bine trees is to arrange ranks in a binomial tree so that communicating ranks are closer than in standard binomial trees. Reducing this distance increases the likelihood that communication remains within the same group, thereby benefiting from higher bandwidth and lower latency. To keep the algorithm simple and generic, and not tied to any specific topology, standard binomial trees do not use the actual physical distance between nodes (e.g., obtained from their location (6061055)); instead, they determine distances based on the difference in rank identifiers.
@@
For example, in distance-halving binomial trees, rank $6$ first communicates with rank $4$ (distance $2$) and then with rank $7$ (distance $1$). This scheme reasonably approximates physical distance when ranks are mapped linearly across the system, which is often the case (e.g., under Slurm’s default block distribution (slurm)). If this is not the case, a remapping to a block distribution could be applied when the communicator is created, and the remapped rank identifiers used instead of the original ones. On the systems used in this work, for instance, hostnames are numbered consecutively across groups or subtrees, so it is sufficient to sort ranks by hostname.
Since our goal is to design an algorithm as simple and generic as standard binomial tree algorithms, we adopt a similar approach but use the modulo distance between rank identifiers. Namely, we consider the $p$ ranks to be arranged in a circle $0,1,\ldots,p-1$ and define the distance between two ranks $r$ and $q$ as the minimum distance along the circle, i.e., $d(r,q)=\min((r-q)\bmod p,\,(q-r)\bmod p)$. To motivate this choice, consider three ranks ($0,1,2$), each in a different group. Standard binomial trees would regard the distance between ranks 0 and $2$ as larger than that between ranks 0 and $1$, whereas in practice, both involve inter-group communication and should therefore be treated as equal. Modular distance achieves this. Of course, counterexamples exist where the standard distance is more accurate. However, as shown in Secs. 2.4.2 and 5, for jobs spanning multiple groups, as is often the case, modular distance provides a better approximation of the actual distance, enabling Bine trees to outperform standard binomial trees in most scenarios.
@@
We construct a distance-halving Bine tree recursively. A two-node Bine tree (Fig. 3, order 1) is identical to a standard binomial tree. To build a four-node (i.e., order 2) Bine tree, we connect an order 1 Bine tree (Fig. 3, [Uncaptioned image]), to a mirrored order 1 Bine tree (i.e., reflected on the vertical axis) and placed to its left (Fig. 3, [Uncaptioned image]). The mirrored tree is placed to the left rather than to the right to keep a shorter modulo distance between the roots of the subtrees (one rather than two as in a binomial tree – see Fig. 2, [Uncaptioned image]). Ranks are numbered left to right starting from the root, wrapping around once the rightmost node of the tree is reached (e.g., on the 4-node tree, ranks $2$ and $3$ are those on the left of the root).
Figure 3. Distance-halving Bine trees construction.:
@@
The process generalizes recursively: an eight-node Bine tree is constructed by combining an order-2 Bine tree (Fig. 3, [Uncaptioned image]) and a mirrored order-2 Bine tree (Fig. 3, [Uncaptioned image]), placed side by side and arranged so that the modulo distance between the roots is as small as possible. In this case, placing the new four-node subtree on the right (rather than on the left) yields a root-to-root modulo distance of three, compared to five in the alternative placement. For comparison, the corresponding binomial tree would result in a modulo distance of four (Fig. 2, [Uncaptioned image]). Appendix C discusses how to handle cases where the number of nodes is not a power of $2$. Moreover, if the root is rank $t\neq 0$, we apply a logical rotation by subtracting $t$ from all rank identifiers modulo $p$.
@@ 2.3.2. Determining the Communication Partner
In the broadcast, each rank must determine when to join the tree, or in other words, at which step it will receive data from its parent. We denote by $u$ the number of consecutive least significant bits that are equal to each other, starting from the least significant bit. E.g., for a 16-node Bine tree, $u=3$ for $1000$, and $u=2$ for $1011$. In a distance-halving Bine tree broadcast, rank $r$ receives data at step i=sui=s-u (with $i\in[0,s-1]$). As illustrated in Fig. 4 [Uncaptioned image], for rank $8$, we have $rank2nb(8)=1000$, so it receives data at step i=su=43=1i=s-u=4-3=1.
Figure 4. A 16-node (order 4) distance-halving Bine tree.:
@@
with $\oplus$ denoting the XOR operation. In other words, at step $i$, each rank $r$ that has already received the data will send it to a rank $q$ whose negabinary representation differs in the least significant sis-i bits. For example, as shown in Fig. 4 [Uncaptioned image], at step $i=2$, rank $8$ sends data at rank $7$ since their negabinary representations ($1000$ and $1011$, respectively), differ for the least significant si=42s-i=4-2 bits. For clarity, the complete MPI code of the distance-halving Bine tree algorithm for the broadcast is provided in Listing 1. Notably, the implementation is no more complex than that of standard binomial trees (mpich-bcast-binomial; ompi-bcast-binomial). Appendix C discusses how to handle cases where the number of nodes is not a power of $2$.
@@ 2.4.1. Theoretical Bound Computation
In binomial trees, the maximum distance between ranks is $p/2$, so the distance between communicating ranks coincides with their modular distance. This allows us to compare Bine and standard binomial trees in terms of modular distance and to quantify the reduction in distance between communicating ranks using the ratio:
@@
This implies that in Bine trees, communicating ranks are at a $\sim 33\%$ shorter modular distance compared to standard binomial trees, reducing the probability that communicating ranks are in different groups and communicate through global links.
Quantifying the exact reduction in inter-group traffic is challenging, as it heavily depends on the specific topology and on how ranks are allocated to groups. However, we can estimate an upper bound by considering a worst-case scenario in which each group only has a single outgoing link to another group. Let us focus on a specific step $i$ and assume the group size equals $\delta_{binomial}(i)$, so that all communications in the standard binomial tree cross group boundaries. In this worst-case scenario, all $\delta_{binomial}(i)$ communications are directed to nodes in other groups, whereas for Bine trees this only happens for $\delta_{bine}(i)$ communications. Thus, the maximum reduction in global link traffic corresponds to the ratio $\delta_{bine}(i)/\delta_{binomial}(i)$, which, as shown in Eq. 2, is equal to $2/3$. In other words, Bine reduces by at most $33\%$ the traffic on global links.
@@ 2.4.2. Empirical Analysis
We additionally validated our assumptions empirically by analyzing one-week job allocation data from the Leonardo (turisini2023leonardo) supercomputer and two-weeks data from the LUMI (lumisupercomputerLUMIsFull) supercomputer. Using Slurm commands (squeue and scontrol), we obtained the list of jobs running on the system and, for each job, the hostnames of the allocated nodes. We then identified the Dragonfly/Dragonfly+ group to which each node was assigned. On Leonardo, we relied on existing hostname-to-group mappings (leonardomap), while on LUMI we extracted this information using the sinfo -N -o "%N %f" command. With this data, we determined the group assignment of each rank for every job. Based on this mapping we estimated, for each job, the global traffic of an allreduce operation using standard binomial and Bine trees, and computed the resulting traffic reduction achieved by Bine across allocations.
Figure 5. Reduction in global traffic on Leonardo and LUMI.:
We present this analysis in Fig. 5, where we report the distribution of global traffic reduction across jobs using boxplots, grouped by the number of nodes. The data includes a total of $1116$ jobs on Leonardo and $1914$ jobs on LUMI. The largest observed jobs spanned up to $256$ nodes on Leonardo and $2048$ nodes on LUMI, matching the maximum job sizes allowed on the two systems. For clarity, outliers are not shown in the figure. No outliers above the theoretical upper bound of 33% were observed, coherently with our model (Eq. 2). In a few cases, all occurring on less than 64 nodes, Bine trees increase global traffic. As discussed in Sec.2.2, this happens because Bine minimizes distance modulo $p$ rather than the actual distance. In contrast, for larger allocations, Bine consistently outperforms standard binomial trees. The reduction in global link traffic increases with the number of nodes, as larger jobs typically span more groups, creating more opportunities for Bine to reduce inter-group communication.
@@ 3.1. Bine Butterflies
In a butterfly communication pattern, each rank exchanges data with another rank at every step. A Bine distance-halving butterfly can be seen as the superposition of multiple distance-halving Bine trees, each rooted at a different rank. The left part of Figure 6 illustrates this view, highlighting the trees rooted at ranks 2 and 3 with solid orange and dotted red lines, respectively.
Figure 6. An 8-node distance-halving Bine butterfly (left), and distance-doubling Bine butterfly (right).:
To guarantee correctness, these trees must be arranged so that the send and receive decisions are consistent across all ranks. We start from the distance-halving tree rooted in rank 0. As discussed in Sec. 2.3, at step $i$ rank 0 communicates with a rank whose negabinary representation has the least significant sis-i bits set to 1, and the remaining bits set to 0, i.e., with rank:
@@
Even ranks $r$ communicate as rank 0, but rotated to the right by $r$ positions. Moreover, since each rank communicates with another one whose negabinary representation differs in the last $i+1$ bits, even ranks always communicate with odd ranks, and vice versa. Consequently, odd ranks $r$ communicate in a mirrored way with respect to the even rank r1r-1. Thus, the communication partner $q$ for a rank $r$ at step $i$ can be determined as:
@@
Since the distance between communicating ranks in Bine butterflies matches that of Bine trees, distance-halving Bine butterflies reduce global traffic by up to $33\%$ compared to standard distance-halving butterflies. Distance-halving butterflies are advantageous for collectives that increase the size of the exchanged data at each step, such as allgather, since in the last steps, where more data is transmitted, communication happens between closer nodes. Alternatively, it is possible to build a distance-doubling Bine butterfly by inverting the communication order, as shown in the right part of Fig. 6. This is preferable for collectives that exchange more data in the first steps, such as reduce-scatter (Sec. 4.3). Appendix C discusses how to handle cases where the number of nodes is not a power of $2$.